flying with the wind, a jet can travel the 4200 km distaance between san francisco and new work in 6 hour. the return trip against the wind takes 7 hours. find the rate of the jet in still air and the rate of the wind
Answer by mananth(16949)

plane speed=x mph
wind speed=y mph
with wind Speed = (x+y)
against wind speed= (x-y)
d= 4200
4200 / (x+y= 6
divide by 6
700 / (x+y)= 1
x+ y 700 ............1
d= 4200
4200 /(x-y)= 7
divide by 7
600 /(x-y) = 7
x-y = 600.00 .............2
add up (1) & (2)
2 x = 1300
/ 2
x= 650 mph speed of plane in still air
plug value of x in (1)
we get y= 50 mph speed of wind